Daily Reports

Because of COVID-19, only IRAM staff is allowed to observe on-site until new notice.

Back to the NIKA2 Pool page

Nov 2nd, Tuesday

21:00 - 03:00 Angel, Ioannis: After weekly maintenance, telescope was parked due to bad weather conditions. Very high tau and variations fog building up around the telescope.

Nov 3rd, Wednesday

03:00 - 09:00 Ioannis: The bad weather conditions continued throughout the night with fog, strong winds and icy rain (!). Operator turned on the deicing mechanism to make sure that we can get back to observing conditions faster as the weather improves. [Note that it will take several hours for the telescope to relax after switching off de-icing. (CK)] Conditions are expected to improve later in the morning..

Some snapshots from this morning:
attachment:IMG_9373.jpg attachment:IMG_9374.jpg

09:00 - 15:00 Ioannis: Around 09:30 (UT) the weather conditions improved (tau ~0.26) and we started with the NIKA2 observations for the 2nd pool week. With the telescope at horizontal coordinates [140,40], the NIKA2 acquisition was initiated at the usual frequency (~24 Hz). The initial tuning of the resonances resulted in bad pixel numbers of 41/5/42 for Ar1/Ar2/Ar3, respectively. The noise traces were inspected and found to be very steep and dominated by the low frequencies, possibly due to the still unstable weather. No feature was found around 1 Hz, so there doesn't seem to be any issue with the PT. Then Mercury was used for initial pointing and focus. The baseline fluctuations were about 6-15 Jy/beam (rms) @ 2mm and about 30 Jy/beam (rms) @ 1mm. Beam size was 17.4" x 18.8" @ 2mm. With relatively stable baselines, a calib_1scan was performed on Mercury (scan 20211103s14) and a skydip (scan 20211103s11). Afterwards, start of observations for the DDT project E03-20. After pointing on 1226+023, we did 2 repetitions of the OTFmaps of the target source. Around 13:00 (UT) the weather conditions deteriorated, the beam broadened to about 19" @ 2mm and 13-14" @ 1mm and the baseline variations significantly increased (+/- 20 Jy/beam @ 2mm and +/- 40 Jy/beam at 1mm). Therefore, we continued with the time filler project 045-21 (during tau variations 0.3 - 0.5) until 14:30 (UT). At 13:55 (UT), "level 6: attente" was selected at the cryostat window, it was at "level 1: auto" until then.

15:00 - 21:00 Isabelle and Karine: Tau range is 0.2 - 0.3 at the beginning of the shift. The noise trace behavior is generally similar as the previous shift and the number of bad pixels are 58/4/58 for Ar1/Ar2/Ar3, respectively. After restarting the NIKA2 acquisition at ~48 Hz, observations for project 086-21 (ARSENAL) start. The opacity was better than the predictions. We did 2 repetitions on A2163 and 1 on A2218. Tau was < 0.3. Then we tried to do a skydip but the telescope control system failed. We experienced several stuck scans and all measurements in the scan range 79-89 should be discarded. The control system problem was resolved at 19:00 (UT) (~1hr lost), then focus and pointing. We switched to project 035-21 and observed 1 repetition on B335.

The sky is finally getting clearer:

21:00 - 03:00 Philippe and Michael: Very good weather conditions (stable tau ~ 0.2 throughout the shift). Obtained several maps for project 018-21. Then did some calibrations (2 skydips [20211104s5 and 20211104s7], 1 calib_1scan [20211104s3] + 1beammap1scan99sub h [20211104s4] on Uranus) when the sources of 018-21 were too high in elevation. Finally, switched to project 007-21 and launched one set of maps (@mapping_1390_v2) for this project at the end of shift. [Had to re-launch the script after the scan got stuck once]

Nov 4th, Thursday

03:00 - 09:00 Golshan: We were blessed with very stable weather, with tau varying around 0.2 and rms noise of array 2 in range 3-5 Jy/beam. After a calib1_scan [20211104s14] and beammap [20211104s15] on Uranus, I started observing NGC3521 and I could finish 13 repetitions of its script. I renewed the pointing every hour and corrected the focus once in the beginning and once in the middle of the shift. May the good weather continue! [It was reported that PIIC monitor for arrays 1 and 3 got stuck a couple of times, possibly due to a faulty scan (header?). The problem was resolved by restarting the monitors. (IM)]

09:00-15:00 Yali, Ioannis: Many thanks for the nice training for the observing with NIKA2 by Ioannis at the beginning of the shift. After the pointing and focus on Mercury, we did calib_1scan with a scan number of 20211104s69, and after which, we did a scan of skydip [20211104s70]. The weather was very good and stable with Tau225 of ~0.2. Then we moved to the project 083-21. We observed two targets (J1530+1049 and J1110+6102) of project 083-21 with very stable sky conditions (baseline variations ~0.5 Jy/beam at 2mm). During the last few scan on target J1110+6102, the weather became a little bit unstable (clouds around the telescope; baseline variations > 20 Jy/beam at 2mm). So after the observations on target J1110+6102, we decided to change to project 045-21. Before the project 045-21, we corrected the pointing and focus on 1226+023. The Tau225 was steadily increasing up to ~0.35-0.4. Then we observed targets in the fields GP005_XXX and GP006_XXX.

Ioannis: It was reported and confirmed that array 1 shows a persistent feature in the lower left part of the noise distribution (see plot attached below) with noise levels x2 or x3 compared to the rest of the array. I tried to retune NIKA2 for a few times but the feature didn't go away. It appears in all scans of array 1 at the same position. Any ideas/suggestions are most welcome.


15:00 - 21:00 Karine and Isabelle: At the beginning of the shift project 045-21 was being observed. Tau increased during the shift to an average of 0.4 with peaks at 0.7. We continued to observe project 045-21 up to 17:30. Then during sunset it was hard to get a good focus and we had to make several tries. We did a skydip (20211104s134). Before being able to get a good focus it started to rain, around 18:45. The weather finally got much better around 20:00 and we started observing again at 20:45 project 035-21 and did one rep on B335 with tau < 0.1!

21:00 - 03:00 Philippe and Michael: Fairly good weather conditions (tau_225 between ~0.1 and ~0.3, with some fluctuations during the shift); no technical problem. Observed project 035-21 (source B335) for 1 hour. Then switched to project 018-21 for ~ 2.5 hour (6 maps) until the Perseus sources of this project reached an elevation > 75 deg. Subsequently observed sources B213 and B211 of project 122-16 (GASTON) for ~ 1.3 hour until they also reached a very high elevation. Then did some calibration observations on Uranus: 1 calib_1scan [20211105s10] + 1 beammap1scan99sub [20211105s11]. Ended with a skydip [20211105s17] (overlap with the next shift).

Nov 5th, Friday

03:00 - 09:00 Golshan: The weather was well throughout my shift. Apart from a few spikes, tau kept decreasing gradually (converging to ~0.15), so I could carry on with observing IMEGIN galaxies. After 2 repetitions on NGC628, which was going to set, I did a skydip (20211105s17, left from the previous shift). Then I completed 5 repetitions on NGC3521 and 8 repetitions on NGC4321 to the end of the shift. I repeated the pointing every hour and corrected the focus every ~2.5 hours. In the end, I did a calib_1scan on Mercury and another skydip (see scans in following report block) before handing it over to the AoD.

09:00 - 15:00 Ioannis: The weather conditions were very good for the first half of the shift (tau225 ~0.15) and there was some slight increase in tau with small spikes up to ~0.35 in the second half of the shift. The spikes were most likely caused by small clouds passing in the direction of the taumeter. In general the sky was very clear and only distant clouds were sometimes visible at low elevations (<30º). At the transition from the previous shift, I did a skydip (20211105s62) and then pointing and focus on Mercury along with a calib_1scan (20211105s72). Baseline variations at <2 Jy/beam at 2mm. Then I observed project 083-21 (J0730+4811) for about 1 hour. After another pointing and focus on Mercury, I started observing project 084-21 for about 2 hours (2 repetitions on AK4-224). During the next pointing and focus on Mercury, the weather conditions slightly degraded (baseline variations ~3 Jy/beam @ 2mm and ~7-8 Jy/beam @ 1mm). Then a calib_1scan on Mercury (20211105s122) and another skydip (20211105s124). Afterwards the weather became a bit unstable (small clouds around the telescope), so I changed to project 045-21 and observed the field GP027_XXX. At the end of the shift the tau225 again stabilized below 0.2

15:00 - 21:00 Karine and Isabelle: Very good shift! Project 045-21 was being observed at the beginning of our shift (GP027_XXX) until 15:55 (overlap from previous shift). Tau was good and stable (< 0.2) so we switched to project 035-21 (3 repetitions on B335). At 18:00 we did a skydip [20211105s146] and focus (very good) on the nikas-21 account. The pointing were very good too (17x17”). We then continued observing project 035-21 (4 repetitions on B335). The opacity was very good (~0.16) and stable (<0.2) up to 20:00 and with spikes < 0.25 for the last hour.

21:00 - 03:00 Miren, Chenlin and Frédéric: In the first part of the night there were spikes in the tau values (0.2 to 0.4). We have done a beammap (20111105s188)+ calib (20111105s187) on Uranus. Then, we have done 2 repetitions on ACTJ2302 of 199-16. We have done a skydip (20111105s199). Then move to 007-21 … but we encounter a snow episode! The antenna was parked and the shift is over.

Nov 6th, Saturday

03:00 - 09:00 Michael: Observations started again at 5am and the weather was stable around tau 0.2. The first pointing and focus on Uranus (scan 1-6) were accidentally done as project nias-21, so it doesn't appear in the nikas-21 account of TAPAS. I then observed project 018-21 for about 1.5h until the sources were setting. Following focus on 0316+413 that was setting, had differing results for IDL and PIIC and was repeated on the rising source 1226+023 with better results. I then observed project 160-16 for 1h. Then another Focus on 1226+023 for sunrise followed by a skydip [20211106s38]. I then returned to project 160-16 for the rest of the shift.

09:00 - 15:00 Julien and Chenlin: We start from doing calib_1scan toward Mercury [20211106s49]. We move to observing project 160-16 (NGC4321, 1 repetition, 2 scans). The tau225 value is ~ 0.25. The receiver engineer does some tests from 10:20 - 10:30 (local time). Then we continue to observing project 160-16 (NGC4321, 4 repetition, 8 scans). While the tau225 value is increasing with some spikes up to 0.58. At 12:30, as the opacity increases (>0.3), we switch to project 045-21 (GP075_0+0_2, GP075_3-0_2, GP075_3+0_0, GP063_0+0_2, GP063_3-0_2, GP063_3+0_0). We finish by doing pointing and focus toward NGC7027, and doing skydip [20211106s85].

15:00 - 21:00 Karine and Isabelle: We started our shift continuing project 045-21 (part of GP063, GP038). Tau was 0.2-0.3 with spikes up to 0.45 and with +/- 10-20 Jy/beam. We did a calib_1scan on NGC7027 [20211106s98] then a focus and pointing at 18:00. Then we switched to project 035-21 and observed B335 (tau ~0.15, good stability) for one hour until the wind was too strong and the antenna was shaking (1 repetition + 1 repetition done at 3/4 the last scan was cancelled). We tried to observe at lower elevation project 122-16_GP but we stoped during the first scan because of the wind. The operator stopped the telescope at 19:30. We started again to observed at 20:30 on project 035-21 (one rep on B335).

21:00 - 03:00 Chenlin, Xavier and Emmanuel: The weather is excellent with tau225 values ranges from 0.1 to 0.15. We start from doing calib_1scan [20211106s131], beam map [20211106s132] and skydip [20211106s133] toward Uranus. Then we move to project 199-16 (ACTJ2302, 1 repetitions 4 scans, PSZ2G155 4 repetitions 16 scans (completed)). At 01:10, we switch to project 007-21 with doing the pointing of Uranus. The weather is good and stable with tau225 around 0.15. We observe 869 of 007-21 (2 repetitions 8 scans). Tau225 value increases with some spikes up to 0.32 from 02:20. We end this shift by doing pointing and focus toward Uranus. Something wrong with the telescope during doing the focus at 2:35 [stuck scan]. We re-do focus from 2:40.

Nov 7th, Sunday

03:00 - 09:00 Michael: The weather was exceptional stable during the shift, tau 0.14 to 0.12. I started with one script for 007-21, then changed to 018-21 for one field. Then calibration [20211107s34] and beammap [20211107s35] on Uranus before it was setting, and a skydip [20211107s36]. Then I continued with 018-21 for two more fields. For the rest of the shift I switched to 085-21 and did 3 repetitions on the source.

09:00 - 15:00 Chenlin and Julien: The weather is excellent with tau225 values ~0.15. We start from project 052-21 with pointing toward 0736+017. We stop after 2 scans (3 scans for one repetition) since the elevation is too low (<15 deg). 10:15 Some tests by Ioannis. 10:35 Pointing and focus toward Mercury. Then calib_1scan [20211107s85] and skydip [20211107s86], and beam map on Mars [20211107s88]. A spike of tau225 value up to 1 from 11:20 - 12:50. From 11:43, we observe project E03-20 till the end of the shift (3 repetitions, 33 scans). The tau225 value remains stable between 0.13 and 0.15 after 13:00.

15:00 - 21:00 Xavier: Opacity was excellent (0.1) during the whole shift. The observations did not start well with Mars and Mercury completely out of focus. But this was because they are too close to the Sun. Going to NGC7027 and MWC349 improved things but there was a lot of anomalous refraction for a while. We did Gaston LP (one run on each of the 3 GP positions) for 1h40min. We got stuck with the telescope because the focusing source NGC7027 was too high in elevations, then after focusing on 3C434.3 (11.5" & 17.6", 3.9 Jy & 5.2Jy) we did 4 repeats on ACTJ2302 (199-16 LPSZ program). Very smooth run at the end. [Calibrations in this run: 20211107s138 calib_1scan on Mars but maybe not good because of extended beam, 20211107s149 calib_1scan on NGC7027 a bit better, 20211107s152 calib_1scan on MWC349 good, 20211107s153 skydip, 20211107s169 calib_1scan on 3C454.3 (IM)]

21:00 - 03:00 Antonio and Marco: We start with good opacity (tau225) close to 0.1. We begin with project 199-16 [ACTJ2302, 2 repetitions, 4 scans/each 0.10<tau_225<0.11 (completed)), ACTJ2130, 2 repetitions (2&3/37), 4 scans/each tau_225=0.11, stable and better than predictions!]. We move on project nikas-21 to do calib_1scan [20211107s220] towards Uranus and beam map but we encounter problems with FWHMs of 1mm beams (close to 13" with triangle shape "Y") probably due to the high elevation as suggested by AoD and operator, therefore we postpone beam map on Uranus when it would be at a lower elevation. We move again to project 199-16 [PSZ2-G111.6, 1 repetition, 4 scans/each, the opacity is has a small increase 0.11<tau<0.14 (completed)]. Due to problems of large FWHM @1mm (19"!!) towards W3OH, even after focusing, we can't go ahead to observe PSZ2G141, therefore we move back to Uranus for beam map [20211108s13] and then skydip [20211108s14]. The opacity increased during the second half of the shift from 0.10 to 0.17, with spikes.

Nov 8th, Monday

03:00 - 09:00 Chenlin: The tau225 value is about 0.2 and stable (with baseline variation ~ +/- 4 Jy/beam @2mm) at the beginning of the shift. We start by observing source 1390 of 007-21 (4 repetitions, 16 scans). [05:08] we move to Uranus to do pointing, focus, calib_1scan [20211108s44], skydip [20211108s45] and beam map [20211108s46]. [06:08] we observe source 4149 of 007-21 (2 repetitions, 8 scans). The tau225 value rise a bit 0.24, but the weather is still stable with baseline variation +/- 5 Jy/beam @2mm. [07:08] We do pointing and focus toward 0316+413. The tau225 value rise to ~0.26. [07:30] We continue to observe 4149 of 007-21 till the end of the shift (3 repetitions, 12 scans). During the last repetitions, tau225 value increases higher than 0.3, and the baseline variation ~ +/- 10 Jy/beam @2mm

09:00 - 15:00 Ioannis: The tau225 was increasing at the beginning of the shift from 0.25 to 0.35 but the 2mm baseline variations remain consistently around 5 Jy/beam. I start with pointing and focus on Mercury. While waiting for opacity to stabilize (currently at ~0.3), I do a calib_1scan on Mercury [20211108s82], a skydip [20211108s83] and I retune the resonances (dead kids 57/4/44 for A1/A2/A3, respectively). I start with project 160-16 and I do 1 repetition on NGC4536 and 2 repetitions on NGC4321. At that moment, the baseline variations increased further (7-8 Jy/beam @2mm) and the opacity is now above 0.35. At 12:00 I changed to EMIR back up project E04-20 and stayed until the end of the shift (opacity peaked at 0.5).

15:00 - 21:00 Golshan: After a short training session for EMIR observations, the whole shift was devoted to observations of the EMIR backup project E04-20.

21:00 - 03:00 Antonio, Marco and Juan (on site): We start with opacity (tau225) close to 0.3, slowly decreasing. After pointing/calibrations/focusing we move to project 010-21 having tau225 around 0.25 but with fluctuations. Beam map on Uranus (20211108s252) and focus. With the opacity more stable, around 0.22-0.23, we move to 199-16 project [PSZ2G141, 2 repetitions (2/4), 4 scans/each], pointing/focus on Uranus and the back on [PSZ2G141, 2 repetitions (4/4), 4 scans/each (completed)].

Nov 9th, Tuesday

03:00 - 09:00 Chenlin: The tau225 value is about 0.25 (with baseline variation ~ +/- 4 Jy/beam @2mm) at the beginning of the shift. We start by doing pointing and focus and skydip [20211109s32] toward Uranus. Beammap had been done at 00:30. Then we observe source 4149 of 007-21 (3 repetitions, 12 scans). The tau value is stable, from 0.21 to 0.23. [06:00] We move to 0316+413 to do pointing, focus. [06:35] We observe source 1390 of 007-21 (4 repetitions, 16 scans). The tau225 value is about 0.24, but baseline variation ~ +/- 10 Jy/beam @2mm. We do pointing, focus and calib_1scan [20211109s82] toward Mercury at the end of the shift.

09:00 - 10:30 Ioannis: After pointing and focus on 1226+023, I start with project 160-16 and do 3 repetitions (6 scans) on NGC4321 (completed).

DailyReportsNika2Pool021121 (last edited 2021-11-09 09:35:00 by NikaBolometer)